OpenStudy (anonymous):

Can SOMEONE help me for a medal and fan???????????????????????????????????!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Find the slope of the line that passes through the point (3, -9) and is parallel to y = 7x + 8. The slope is ___.

OpenStudy (misty1212):

slope of \(y=7x-8\) is \(7\) right?

OpenStudy (misty1212):

use the point slope formula and get \[y+9=7(x-3)\]

OpenStudy (anonymous):

so the slope is y+9=7(x-3)? @misty1212

OpenStudy (anonymous):

If you're given an equation of a line, the slope is given (the number in front of the X). Parallel lines have the same slope, so the slope is still 7 :) Hope this helps!

OpenStudy (anonymous):

Parallel lines have the same slope. the slope of the equation give is 7 so the paralle line is also 7